3.011 x 10^22 atoms of an element weight 1.15 gm. What is the atomic mass of the element?
1
Expert's answer
2020-11-22T11:11:28-0500

At first, number of moles should be found:

n=N/Na

n=3.011×1022/6.022×1023=0.05(mol)

M=m/n=1.15/0.05=23(g/mol)

So, we may assume that this element is Sodium.